# B2N3D: Progressive Learning from Binary to N-ary Relationships for 3D Object Grounding

### v1

@@ IV-C Comparison with State-of-the-Art Methods
To evaluate the effectiveness of our proposed B2N3D, we compare it with state-of-the-art approaches on Nr3D, Sr3D, and ScanRefer datasets. As shown in Table I, our method achieves superior performance in overall localization accuracy. Specifically, B2N3D achieves 62.7% on hard samples and 67.8% on view-dependent samples of Nr3D, which are significantly higher than the latest method xM_Match, demonstrating our advantage in challenging cross-modal understanding. The synthetic dataset, Sr3D, has a lower proportion of complex relation descriptions (Figure 7). Thus, the enhancements brought about by our progressive learning method are subtly noticeable.
@@
In addition, the qualitative results in Figure 5 further validate the effectiveness of our method in precisely localizing objects defined in natural language descriptions. This is particularly notable in complex scenarios where the target location hinges on multiple referential relations. Overall, these results demonstrate that our method not only achieves competitive performance in 3D object grounding but also offers global perception advantages in terms of complex relationships.
@@ IV-D1 Components of Relational Learning
As (a) in Figure 5, it is necessary to combine the multiple positional relationships between “chair” and “wall”, “table”, and “door” to select the correct “chair” from the four. Our proposed B2N-PRL module enables precise localization of the target through global perception. For targets that can be identified based on simple contextual cues, such as the “backpack” in (c), where its spatial relationship with either the “floor” or the “whiteboard” is sufficient for localization, correct results can still be achieved without global relational learning.
@@ IV-E Discussion about Soft Relational Labels
Figure 7 illustrates the perception of relationship numbers in soft labels across various datasets. Both Nr3D and ScanRefer consist of human-annotated descriptions with more complex referential expressions, resulting in a higher number of binary relationships involving more than two entities. In contrast, Sr3D is a template-based synthetic dataset, where single spatial relations dominate. As shown in the statistical results, multi-relational descriptions occupy a considerable proportion in the datasets, indicating that learning from single binary relationships alone is insufficient for understanding complex textual inputs. Experimental results shown in Table VI demonstrate that our proposed B2N-PRL module significantly improves localization performance for understanding multiple relations. The largest gain is seen on the Nr3D dataset, where accuracy increases from 60.6% to 67.7%.

### v2

@@ IV-C Comparison with State-of-the-Art Methods
To evaluate the effectiveness of our proposed B2N3D, we compare it with state-of-the-art approaches on Nr3D, Sr3D, and ScanRefer datasets. As shown in Table I, our method achieves superior performance in overall localization accuracy. Specifically, B2N3D achieves 62.7% on hard samples and 67.8% on view-dependent samples of Nr3D, which are significantly higher than the latest method xM_Match, demonstrating our advantage in challenging cross-modal understanding. The synthetic dataset, Sr3D, has a lower proportion of complex relation descriptions (Fig. 7). Thus, the enhancements brought about by our progressive learning method are subtly noticeable.
@@
In addition, the qualitative results in Fig. 5 further validate the effectiveness of our method in precisely localizing objects defined in natural language descriptions. This is particularly notable in complex scenarios where the target location hinges on multiple referential relations. Overall, these results demonstrate that our method not only achieves competitive performance in 3D object grounding but also offers global perception advantages in terms of complex relationships.
@@ IV-D1 Components of Relational Learning
As (a) in Fig. 5, it is necessary to combine the multiple positional relationships between “chair” and “wall”, “table”, and “door” to select the correct “chair” from the four. Our proposed B2N-PRL module enables precise localization of the target through global perception. For targets that can be identified based on simple contextual cues, such as the “backpack” in (c), where its spatial relationship with either the “floor” or the “whiteboard” is sufficient for localization, correct results can still be achieved without global relational learning.
@@ IV-E Discussion about Soft Relational Labels
Fig. 7 illustrates the perception of relationship numbers in soft labels across various datasets. Both Nr3D and ScanRefer consist of human-annotated descriptions with more complex referential expressions, resulting in a higher number of binary relationships involving more than two entities. In contrast, Sr3D is a template-based synthetic dataset, where single spatial relations dominate. As shown in the statistical results, multi-relational descriptions occupy a considerable proportion in the datasets, indicating that learning from single binary relationships alone is insufficient for understanding complex textual inputs. Experimental results shown in Table VI demonstrate that our proposed B2N-PRL module significantly improves localization performance for understanding two or more referential relationships. The largest gain is seen on the Nr3D dataset, where accuracy increases from 60.6% to 67.7%.